OpenStudy (anonymous):

two buses leave dallas at the same time and travel in opposite directions. bus A averages 58mph and bus B averages 52mph when will they be 363 miles apart? please also tell me how you set up the equation

OpenStudy (anonymous):

r t = d Solve the following for t, the time. 58 t + 52 t = 363 t = 33/10 = 3.3 = 3hrs 20 minutes after they start.

OpenStudy (anonymous):

ohh ok thanks so much!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

do you want to help me with another one? lol

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

there is no closure property of division that applies to integers. ex: 2 divided by 3 is not an integer. what is another example of a set of real numbers that does not have a closure property for one of its basic operations? give an example to illustrate your claim.

OpenStudy (anonymous):

I am guessing on this one.\[\sqrt{2} \]

OpenStudy (anonymous):

but shouldnt it be {+ - x or/}?

OpenStudy (anonymous):

http://www.abstractmath.org/MM/MMIntegers.htm Refer to the above where they cover +, -, * and /.

OpenStudy (anonymous):

thanks!

OpenStudy (turingtest):

The set they are asking for is then the set if irrational numbers. This set is not closed because\[\sqrt{2}*\sqrt{2}=2\]therefor the product of two irrational numbers may be rational.

OpenStudy (turingtest):

The set of imaginary numbers is not closed either, I'll leave it up to you to show that.